In a coffee-cup calorimeter, 57.0 mL of 0.100 M AgNO3 and 57.0 mL of 0.100 M HCl are mixed to yield the following reaction.
Ag+(aq) + Cl −(aq) → AgCl(s)
The two solutions were initially at 21.90°C, and the final temperature is 22.70°C. Calculate the heat that accompanies this reaction in kJ/mol of AgCl formed. Assume that the combined solution has a mass of 114.0 g and has a specific heat capacity of 4.18 J/°C · g.
1
Expert's answer
2019-11-13T08:29:32-0500

The reaction is expressed with the following equation:


"AgNO_3+HCl \rightarrow AgCl \downarrow+HNO_3;"


Thus, one mole of AgNO3 and one mole of HCl produce together one mole of AgCl, while in our case the amount of AgNO3 is:


"n(AgNO_3)=n(AgCl)=C_{AgNO_3}*V_{AgNO_3}=0.100M*0.0570L=5.7*10^{-3}mol;"


The amount in moles of HCl and AgCl precipitate is the same.

Now, we have to calculate the amount of heat released in the cup calorimeter. Hypothetically, all heat was absorbed by the solution:


"Q=m(solution)*c(water)*(t(final)-t(starting))=114g*4.18 \frac{J}{C*g}*(22.70C-21.90C)=381J;"


Finally, the molar heat of formation for AgCl precipitate is:


"H=\frac{Q}{n(AgCl)}=\frac{381J}{5.7*10^{-3}mol}=66.8\frac{kJ}{mol}."